A butterfly net hangs from a circular loop of diameter 430mm. You hold the loop horizontally in a region where electric field is 149 N/C

What is the electric Flux through the net?

Does Flux increase or decrease if you rotate loop so that it lies vertical in a plane?
1
Expert's answer
2021-09-09T18:13:57-0400

By definition, the flux is given as follows:


"\Phi = ES\cos\theta"

where "E = 149N\/C" is the field (should be constat), "S = \dfrac{\pi d^2}{4}" is the surface area (in this case circle with diameter "d = 430mm = 0.43m"), and "\theta" is the angle (from the normal to the surface) between field and surface. In the first case "\theta = 0\degree", since the net is held horizontaly.

Thus, obtain:


"\Phi = ES\cos\theta = \dfrac{\pi Ed^2}{4} \approx 21.6\dfrac{Nm^2}{C}"

In the second case the angle become "\theta = 90\degree", then "\cos\theta = 0" and the flux is 0.


Answer. 1) "21.6\dfrac{Nm^2}{C}", 2) .